OpenStudy (anonymous):

approximation of tan 28 degrees -_-''

OpenStudy (anonymous):

Well it's close to 30 degrees.

OpenStudy (anonymous):

Well that's VERY close to tan (30 degrees) which is 1/3

OpenStudy (anonymous):

There is less than a 8% diffrerence.

OpenStudy (anonymous):

Yes, I know that, i have to use newtons method though

OpenStudy (anonymous):

Ohh dear!!!

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

Yepp ! its such a pain!

hero (hero):

lol

OpenStudy (anonymous):

newtons method is a good way to approximate though =]

OpenStudy (turingtest):

you mean the \[f(x)\approx f(x_0)+f'(x_0)(x-x_0)\]thing?

OpenStudy (anonymous):

I can't wait to learn toylor polynomials to approxmitate though.

OpenStudy (anonymous):

Yeah that's the one @TuringTest .

OpenStudy (anonymous):

Yes @TuringTest

hero (hero):

Yeah, the thing that you can simply program into your calculator to have it do it for you automatically

OpenStudy (anonymous):

But what if you can't use a calculator? :O .

OpenStudy (turingtest):

fill in the blanks with \(x_0=30^{\circ}=\frac\pi6\)

hero (hero):

Then you have to do it manually, which can really be a pain.

OpenStudy (anonymous):

@Hero I have to do it by hand :/ no calculators! and ok

OpenStudy (turingtest):

first off, what's\[f(x_0)\]in our case?

OpenStudy (anonymous):

tan x

OpenStudy (turingtest):

that is our \(f(x)\) but we are looking at \(f(x)\) near \(x_0=\frac\pi6\) I asked you what \(f(x_0)=f(\frac\pi6)\) is because that is the first thing we see in our formula

OpenStudy (anonymous):

OH thats what the question was. Yes its is pi/6 b/c 30degrees

OpenStudy (turingtest):

yeah, so what is tan(pi/6) is what my question amounts to

OpenStudy (anonymous):

sqrt3/3

OpenStudy (turingtest):

right, so that's one piece of the puzzle now what about f' ?

OpenStudy (anonymous):

1/cos^2(pi/6)

OpenStudy (turingtest):

right, that's f'(pi/6), which is what?

OpenStudy (anonymous):

1/cos(pi/6)= sqrt3/3 right? but what is it ^2??

OpenStudy (turingtest):

cos(pi/6)=(sqrt3)/2

OpenStudy (turingtest):

so 1/cos(pi/6)=2/(sqrt3) what is 2 squared? what is sqrt3 squared?

OpenStudy (anonymous):

4/3?

OpenStudy (turingtest):

yes :) so now you can fill in the blanks to get a linear approximation for f(x) near pi/6

OpenStudy (turingtest):

\[f(x)\approx f(x_0)+f'(x_0)(x-x_0)\]you know each part now

OpenStudy (anonymous):

so i then do rt3/3+ 4/3(pi/90)= my answer?

OpenStudy (turingtest):

you forgot about \(x_0\)

OpenStudy (turingtest):

oh I see what you did, but it should be negative because \(x<x_0\)

OpenStudy (anonymous):

i thought (xo) was sqrt3/3 ?

OpenStudy (turingtest):

yes that is f(x_o), but I am talking about x_o=pi/6

OpenStudy (turingtest):

better to call it 30 deg right now

OpenStudy (anonymous):

And now you have to use newtons method to approximate the square root of 3?

OpenStudy (anonymous):

so pi/6 + sqrt/3 - 4/3(pi/90) ?

OpenStudy (turingtest):

so\[f(x)\approx f(x_0)+f'(x_0)(x-x_0)=\frac{\sqrt3}3+\frac43(28^{\circ}-30^{\circ})\]

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

wait where did that other pi/6 come from?

OpenStudy (anonymous):

you said i was missing the xo pi/6 lol confused me... the degrees..do i have to convert them to radians?

OpenStudy (turingtest):

yes, because radians have no units, you need a unitless quantity for the value of tangent

OpenStudy (anonymous):

so essentially the 28-30 degrees actually is pi/6-pi/90 ?

OpenStudy (turingtest):

pi/6=30 28-30=-2 (degrees) =-pi/90 radians

OpenStudy (anonymous):

sorry, im confused. how do i know that -2 degrees is pi/90 ?

OpenStudy (anonymous):

@jennilalala He converted degrees to radians.

OpenStudy (anonymous):

I think...

OpenStudy (turingtest):

yep -2 degrees is -1/180th of a circle, which is 2pi radians -2pi/180=-pi/90

OpenStudy (turingtest):

in general\[x~\text{radians}\times\frac{180}\pi=y~\text{degrees}\]and vice versa

OpenStudy (turingtest):

\[y~\text{degrees}\times\frac\pi{180}=x~\text{radians}\]

OpenStudy (anonymous):

rt3/3+4/3(-pi/90)= rt3/3-4pi/270 ...??

OpenStudy (turingtest):

yes, that looks right :)

OpenStudy (anonymous):

so I can leave that as the answer? how do you subtract that ive never even done that before

OpenStudy (turingtest):

there isn't much to do, you can't simplify pi or sqrt3. You could get a common denominator if you wanted, but I see little point in that.

OpenStudy (anonymous):

i see. thank you very much for this detailed explanation/walkthrough

OpenStudy (turingtest):

welcome you should also always check your answer against the value your calculator gives you for tan(28) I get 0.5308... for the approximation and 0.5317... for the actual value, so it's pretty darn close :)

OpenStudy (anonymous):

I wish that he could just write tan 30 :/ . It's pretty close.

OpenStudy (anonymous):

@Dido525 hahahaha me too man. me too...

OpenStudy (anonymous):

&you mean *we? as in students...@Dido525

OpenStudy (turingtest):

but it's not nearly as close as our approximation, I think if you used tan(30) in place of tan(28) our satellites would be flying off into space :P

OpenStudy (anonymous):

I said he :P .

OpenStudy (anonymous):

No kiding!!! @TuringTest

OpenStudy (anonymous):

that's happened before actually, something like that where scientist used the wrong approximation and something space completely messed up

OpenStudy (turingtest):

I remember something like that, a guy in ASA not converting feet to meters. I think I was in high school.

OpenStudy (turingtest):

NASA*

OpenStudy (anonymous):

The didn't convert units. So they ended up making a space probe which crashed on mars because they didn't convert units properly.